OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (king):

y=kx where k is the proportionality constant so 8=-6k so,k=-4/3 so, when x=-2 y=-4/3*-2 =8/3

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2? A. y = –0.75x; 1.50 B. y = –1.33x; 2.67 C. y = 1.33x; –2.67 D. y = 0.13x; –0.25